OpenStudy (anonymous):

Undergraduates in a large university class were asked to record the time they spent on the most recent homework. In a histogram of the recorded times, the height of the bar over the interval “1 hour to 3 hours” is 4% per hour. The height of the bar over the interval “3 hours to 6 hours” is 2% per hour. If the histogram is redrawn with just a single bar over the interval “1 hour to 6 hours”, what should the height of the bar be?

OpenStudy (anonymous):

3-1=2*4%=8% 6-3=3*2%=6% 8%+6%=14%/ 5 =2.8 Let me know if it helped!

OpenStudy (anonymous):

birdy14 thanks